X-ray radiation falls in the wavelength region of 1.00×10-11 to 1.00×10-8 meters. What is the frequency of X-ray radiation that has a wavelength of 1.00×10-11 m?

Frequency = sec-1
Answer #1

Answer:- frequency = 3.00 * 1019sec-1

Solution:-

since we know that,

frequency * wavelength = speed of light

frequency = speed of light / wavelength .......(1)

speed of light = 3.00×108 ms-1

wavelength = 1.00×10-11 m

from equation (1)

frequency = 3.00×108 ms-1 / 1.00×10-11 m

frequency = 3.00 * 1019sec-1